### v1

@@ 3 Monte Carlo samples
The IDM signal samples are generated using MADGRAPH5_aMC@NLO v2.8.1 Alwall:2014hca interfaced with PYTHIA v8.2 Sjostrand:2014zea, using $e^{+}e^{-}$ collisions at $\sqrt{s}=$240 and 365 GeV. The input UFO Degrande:2011ua model is taken from Ref. Goudelis:2013uca. Instead of targeting individual pair production modes, two final states are generated directly, therefore taking into account all contributing diagrams and the proper interferences. The two final states we consider here are given by $HH\ell\ell$ and $HH\ell\ell\nu\nu$, with $\ell=e$, $\mu$ or $\tau$, and $H$ the stable dark scalar [^4] Here and in the following, we use the shorthand notation $\ell\ell\,\equiv\,\ell^{+}\,\ell^{-}$ and $\nu\,\nu\,\equiv\,\nu\,\bar{\nu}$. In all processes discussed here, both electroweak charge and lepton number are conserved.. The lepton charges are omitted in the notations. The $\tau$ leptons are decayed by PYTHIA. Typically, the former final state is obtained with a virtual or real $\mathup{{{Z}}}$ boson (Fig. 1 left), and the sensitivity will roughly depend on the mass difference between $H$ and $A$. The latter final state is mediated via a pair of $\mathup{{{W}}}$ bosons (Fig. 1 middle) and the sensitivity will instead depend on the difference in mass between $H$ and the charged scalars. The signal is generated fixing the coupling $\lambda_{2}$ to 0.1, $M_{H}$ between 70 and 115 (180) GeV in steps of 5 GeV for $\sqrt{s}$ =$240\,(365)\,\text{G\text{eV}}$, and $M_{A}-M_{H}$ between 2 GeV and the kinematic limit for on-shell production in steps of 2 to 5 GeV. Three scenarios are considered for setting $M_{H^{\pm}}$ and $\lambda_{345}$, as detailed in section 2, and 500,000 events are generated per point. The cross section values are shown in Figs. 3 and  4 in the simulated $M_{A}-M_{H}$ vs. $M_{H}$ grid plane for $\sqrt{s}=$240 and 365 GeV respectively, on the left for the $HH\ell\ell$ final state and on the right for the $HH\ell\ell\nu\nu$ final state, for the scenario S1. For the number of events generated, the Monte Carlo (MC) integration error for the cross sections is typically smaller than 0.1%. The minimum transverse momentum of the leptons, $p_{\mathrm{T}}^{\ell}$, is set at 0.5 GeV.
Figure 3: Cross section from the MG5_aMC@NLO simulation at $\sqrt{s}=$240 GeV, for the points simulated in the $M_{A}-M_{H}$ vs $M_{H}$ plane in scenario S1. All cross sections are in fb, computed for $p_{\mathrm{T}}^{\ell}>0.5$ GeV and with a numerical integration error smaller than 0.1%. Left: $HH\ell\ell$ final state, right: $HH\ell\ell\nu\nu$ final state. The dashed line shows the kinematic limit for on-shell production.
@@
Almost no difference is found between S2 and S1, which is expected from the range of $M_{H}$ values of interest here, $M_{H}$$>$$M_{h}$$/2$. The largest differences are for S3 in the $HH\ell\ell\nu\nu$ final state, for which $M_{H^{\pm}}$ drives the cross section. The cross section ratios between S3 and S2 are shown in Fig. 5 in the $M_{A}-M_{H}$ vs. $M_{H}$ plane for $\sqrt{s}=$240 GeV (left) and $\sqrt{s}=$365 GeV (right) for the $HH\ell\ell\nu\nu$ final state.
@@
Background processes include diboson production ($\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$ and $\mathup{{{Z}}}\mathup{{{Z}}}$), inclusive $ee$, $\mu\mu$ and $\tau\tau$ production, SM Higgs boson production in association with a $\mathup{{{Z}}}$ boson, and $\mathup{{{t}}}\mathup{{\overline{{\mathup{{{t}}}}}}}$ production at $\sqrt{s}=$365 GeV. Samples were produced centrally with high statistics, within the so-called Winter2023 campaign winter23, as given in Table 1 for $\sqrt{s}=$240 GeV and in Table 2 for $\sqrt{s}=$365 GeV, using PYTHIA v8.306 Bierlich:2022pfr for the diboson and $\mathup{{{t}}}\mathup{{\overline{{\mathup{{{t}}}}}}}$ samples, and WHIZARD v3.0.3 Kilian:2007gr interfaced with PYTHIA v6.4.28 Sjostrand:2006za for the other processes. For the inclusive $ee$ production, a lower limit of 30 GeV is set on the invariant mass of the pair of electrons produced. The dielectron channel is hence included in the analysis only in the parameter space where the background is simulated, namely for $M_{A}-M_{H}$$>30$ GeV.
The current target values for the total integrated luminosities of the $\sqrt{s}=$240 and 365 GeV FCC-ee runs are 10.8 and 2.7  ab-1, respectively janot_2024_nfs96-89q08.
Both signal and background samples are processed through DELPHES v3.5.1pre05 deFavereau:2013fsa with the IDEA Tassielli:2021rjk detector configuration as per the Winter2023 campaign. Events are produced using the KEY4HEP Ganis:2021vgv framework and analysed using the FCCAnalyses package fccana.
@@
$\mathup{{{Z}}}\mathup{{{Z}}}$ 56162093 1.359 41
$\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$ 373375386 16.4385 23
$\mathup{{{Z}}}$($ee$)$h$ 1200000 0.00716 168
$\mathup{{{Z}}}$($\mu\mu$)$h$ 1200000 0.00676 178
$\mathup{{{Z}}}$($\tau\tau$)$h$ 1200000 0.00675 178
$\mathup{{{Z}}}$($\nu\nu$)$h$ 3500000 0.0462 76
$\mathup{{{Z}}}$($\mathrm{q}\bar{\mathrm{q}}$)$h$ 6700000 0.136 49
$ee$ $15030<M_{ee}<$ 150GeV 85400000 8.305 10
@@
Table 1: Background processes generated in the Winter2023 campaign at $\sqrt{s}=$240 GeV. $\mathup{{{Z}}}\mathup{{{Z}}}$ and $\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$ are generated using Pythia8, whilst the others are generated with Whizard+Pythia6. The last column gives the total integrated luminosity simulated.
@@
$\mathup{{{Z}}}\mathup{{{Z}}}$ 11470944 0.6428 18
$\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$ 11754213 10.7165 1.1
$\mathup{{{Z}}}$($ee$)$h$ 1000000 0.00739 135
$\mathup{{{Z}}}$($\mu\mu$)$h$ 1200000 0.004185 287
$\mathup{{{Z}}}$($\tau\tau$)$h$ 1100000 0.004172 264
$\mathup{{{Z}}}$($\nu\nu$)$h$ 2200000 0.05394 41
$\mathup{{{Z}}}$($\mathrm{q}\bar{\mathrm{q}}$)$h$ 2400000 0.032997 73
$ee$ $15030<M_{ee}<$ 150GeV 3000000 1.527 2.0
@@
$\mathup{{{t}}}\mathup{{\overline{{\mathup{{{t}}}}}}}$ 2700000 0.8 3.4
Table 2: Background processes generated in the Winter2023 campaign at $\sqrt{s}=$365 GeV. $\mathup{{{Z}}}\mathup{{{Z}}}$, $\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$ and $\mathup{{{t}}}\mathup{{\overline{{\mathup{{{t}}}}}}}$ are generated using Pythia8, whilst the others are generated with Whizard+Pythia6. The last column gives the total integrated luminosity simulated.
@@ 4 Definition of the objects in Delphes
$\sigma_{E}=\sqrt{E^{2}\times c+E\times s+n}$
@@
Tracking $e$, $\mu$, charged hadrons $p_{\text{T}}$ $>0.1$ GeV, $|\eta|<2.56$ 1
@@
Jets are reclustered from Delphes reconstructed particles after removing the selected electron and muon candidates. The Durham algorithm is used, with exclusive clustering of two jets, and an energy-based scheme Cacciari:2011ma. The missing energy is taken as given by Delphes, where the DM candidates count as invisible particles.
@@ 5 Background rejection
5.1 Selection
To remain fully sensitive to all signal points but significantly reduce the background contributions, the following preselection criteria are applied, which are inspired from previous studies performed at generator level with the CLIC setup Kalinowski:2018kdn,Zarnecki:2020swm.
Only events with exactly two same-flavour leptons $\ell=e$ or $\mu$, $M_{\ell\ell}<120$ GeV, $|p_{z}(\ell\ell)|<70$ GeV are kept. For the $\sqrt{s}=$365 GeV setup, for several benchmark points the signal contribution has $M_{\ell\ell}$ values well above the Z invariant mass. In order to more efficiently reject the backgrounds, the selection is modified to $|p_{z}(\ell\ell)|<140$ GeV, and the selection in $M_{\ell\ell}$ is chosen to be dependent on $p_{z}(\ell\ell)$, with the formula: $M_{\ell\ell}<(-9.0/14.0\times|p_{z}(\ell\ell)|+200)$, with $M_{\ell\ell}$ and $p_{z}(\ell\ell)$ in GeV. As an example, the distribution of $M_{\mu\mu}$ as a function of $p_{z}(\mu\mu)$ is shown for the sum of the major backgrounds and a few signal representative benchmark points from Ref. Kalinowski:2018kdn in Fig. 6, for $\sqrt{s}=$240 (365) GeV on the left (right). A good agreement is found with the results obtained using the CLIC setup.
Figure 6: Dimuon mass $M_{\mu\mu}$ as a function of the dimuon $p_{z}(\mu\mu)$, for the selection of a muon pair with $e^{+}e^{-}$ collisions at $\sqrt{s}=$240 (365) GeV on the left (right). The sum of the background processes is shown in red. Several representative signal benchmark points are shown in green (BP1), blue (BP2), yellow (BP7) and pink (BP9). On the right, also BP21 is shown in cyan. The specification of the BPs can be found in Ref. Kalinowski:2018kdn. The pink lines highlight the selection requirements.
With a negligible impact on the signal efficiency, the $ee\rightarrow\ell\ell,\,\tau\tau$ background is further reduced by 90% if requiring the transverse missing energy $E_{\text{T}}^{miss}$ $>5$ GeV. As no additional object is expected for signal events, a veto is applied on events with any additional electron or muon, any photon $E>5$ GeV or any jet. To further reject background, only events with the leading (subleading) lepton $p_{\text{T}}$ $<80$ (60) GeV are kept for $\sqrt{s}=$240 GeV, $p_{\text{T}}$ $<140$ (80) GeV for $\sqrt{s}=$365 GeV. Finally, events with $p_{\ell\ell}/E_{\ell\ell}>0.1$ are selected. All selection criteria are summarised in Table 4.
@@ 5.1 Selection
Preselection $|p_{z}(\ell\ell)|<70$ GeV $|p_{z}(\ell\ell)|<140$ GeV $\mathup{{{Z}}}\mathup{{{Z}}}$, $ee\rightarrow\ell\ell$
$M_{\ell\ell}<120$ GeV $M_{\ell\ell}<(-9.0/14.0\times|p_{z}(\ell\ell)|+200)$ $\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$, $ee\rightarrow\ell\ell$
$E_{\text{T}}^{miss}$ $>5$ GeV $\mathup{{{Z}}}\mathup{{{Z}}}$, $ee\rightarrow\ell\ell$
Object veto 3rd lepton $E>5$ GeV, jet, photon $E>5$ GeV $\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$, $\mathup{{{Z}}}\mathup{{{Z}}}$, $ee\rightarrow\ell\ell,\,\tau\tau$
Leptons $p_{\text{T}}$ $p_{\text{T}}$ $<80,60$ GeV $p_{\text{T}}$ $<140,80$ GeV $\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$, $\mathup{{{Z}}}\mathup{{{Z}}}$, $ee\rightarrow\tau\tau$
E/p $p_{\ell\ell}/E_{\ell\ell}>0.1$ $ee\rightarrow\ell\ell,\,\tau\tau$
@@
Just from requiring a pair of same-flavour leptons, 16% (4%) of $\mathup{{{Z}}}\mathup{{{Z}}}$ ($\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$) events are selected, further reduced to 5% (2%) after preselection, and 1.5% (1%) after the veto on other objects and final set of selection criteria. For the dilepton production, the preselection achieves 99% (94%) rejection of $ee$ events in the $ee$ channel for $\sqrt{s}=$240 (365) GeV, 99% for the $\mu\mu$ in $\mu\mu$ channels, and 97% for the $\tau\tau$ events in both channels. The object veto and other selection criteria reduce the $ee$, $\mu\mu$ and $\tau\tau$ backgrounds by another factor of about 40%, 70% and 35%, respectively. For the signal, after requiring a pair of same-flavour leptons with $p>5$GeV  only 2% (0.1%) of events are selected at low mass splitting $M_{A}-M_{H}$, increasing to about 65% (6%) at higher mass splitting, for the production of $HH\ell\ell$ ($HH\ell\ell\nu\nu$) in the scenario S1.
@@
Process $ee$ $\sqrt{s}=$240 GeV $\mu\mu$ $\sqrt{s}=$240 GeV $ee$ $\sqrt{s}=$365 GeV $\mu\mu$ $\sqrt{s}=$365 GeV
@@
$\mathup{{{Z}}}\mathup{{{Z}}}$ 9.74e+04 $\pm$ 1.60e+02 1.09e+05 $\pm$ 1.69e+02 9.30e+03 $\pm$ 3.75e+01 1.05e+04 $\pm$ 3.98e+01
$\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{+}}}\mathup{{{\mathup{{{W}}}}}^{\scriptstyle{-}}}$ 8.02e+05 $\pm$ 6.18e+02 8.72e+05 $\pm$ 6.44e+02 8.75e+04 $\pm$ 4.64e+02 9.32e+04 $\pm$ 4.79e+02
$\mathup{{{t}}}\mathup{{\overline{{\mathup{{{t}}}}}}}$ 0 0
$ee$ 6.16e+05 $\pm$ 8.05e+02 0 1.34e+05 $\pm$ 4.30e+02 0
@@
$\mathup{{{Z}}}$($ee$)$h$ 1.11e+02 $\pm$ 2.68e+00 0 1.88e+01 $\pm$ 6.13e-01 9.98e-02 $\pm$ 4.46e-02
$\mathup{{{Z}}}$($\mu\mu$)$h$ 6.09e-02 $\pm$ 6.09e-02 1.33e+02 $\pm$ 2.85e+00 9.42e-03 $\pm$ 9.42e-03 1.74e+01 $\pm$ 4.05e-01
$\mathup{{{Z}}}$($\nu\nu$)$h$ 2.15e+03 $\pm$ 1.75e+01 2.25e+03 $\pm$ 1.79e+01 6.08e+02 $\pm$ 6.34e+00 6.72e+02 $\pm$ 6.67e+00
$\mathup{{{Z}}}$($\tau\tau$)$h$ 1.62e+01 $\pm$ 9.93e-01 1.90e+01 $\pm$ 1.08e+00 2.41e+00 $\pm$ 1.57e-01 2.11e+00 $\pm$ 1.47e-01
$\mathup{{{Z}}}$($\mathup{{{q}}}\mathup{{\overline{{\mathup{{{q}}}}}}}$)$h$ 0 0 0 0
@@
100-104 $HH\ell\ell$ 1.09e+02 $\pm$ 1.59e+00 1.39e+02 $\pm$ 1.80e+00
100-104 $HH\ell\ell\nu\nu$ 1.96e+01 $\pm$ 1.00e+00 1.96e+01 $\pm$ 1.00e+00
120-231 $HH\ell\ell$ 1.25e+02 $\pm$ 3.47e-01 1.41e+02 $\pm$ 3.68e-01
120-231 $HH\ell\ell\nu\nu$ 2.61e-04 $\pm$ 7.59e-06 2.07e-04 $\pm$ 6.77e-06
125-215 $HH\ell\ell$ 3.14e+02 $\pm$ 8.65e-01 3.53e+02 $\pm$ 9.18e-01
125-215 $HH\ell\ell\nu\nu$ 1.23e-02 $\pm$ 1.14e-04 9.92e-03 $\pm$ 1.03e-04
140-200 $HH\ell\ell$ 3.42e+02 $\pm$ 9.29e-01 3.80e+02 $\pm$ 9.79e-01
140-200 $HH\ell\ell\nu\nu$ 1.31e-02 $\pm$ 1.21e-04 1.06e-02 $\pm$ 1.09e-04
160-185 $HH\ell\ell$ 2.19e+02 $\pm$ 6.29e-01 2.39e+02 $\pm$ 6.56e-01
160-185 $HH\ell\ell\nu\nu$ 3.61e-03 $\pm$ 4.81e-05 2.83e-03 $\pm$ 4.26e-05
Table 5: Expected number of events after selection for both channels ($ee$ and $\mu\mu$) normalised to the total integrated luminosity values of the FCC-ee runs at $\sqrt{s}=$240 and 365 GeV, namely 10.8 and 2.7  ab-1. The signal points are all shown for scenario S1, labelled using the simulated values of $M_{H}$ and $M_{A}$ given in GeV in that order, and separately for the two generated final states, $HH\ell\ell$ and $HH\ell\ell\nu\nu$. Different representative mass points are chosen for the different centre-of-mass scenarios. Only statistical uncertainties are included.

### v2

@@ 3 Monte Carlo samples
The IDM signal samples are generated using MADGRAPH5_aMC@NLO v2.8.1 [32] interfaced with PYTHIA v8.2 [33], using $e^{+}e^{-}$ collisions at $\sqrt{s}=$240 and 365 GeV. The input UFO [34] model is taken from Ref. [35]. Instead of targeting individual pair production modes, two final states are generated directly, therefore taking into account all contributing diagrams and the proper interferences. The two final states we consider here are given by $HH\ell\ell$ and $HH\ell\ell\nu\nu$, with $\ell=e$, $\mu$ or $\tau$, and $H$ the stable dark scalar [^4] Here and in the following, we use the shorthand notation $\ell\ell\,\equiv\,\ell^{+}\,\ell^{-}$ and $\nu\,\nu\,\equiv\,\nu\,\bar{\nu}$. In all processes discussed here, both electroweak charge and lepton number are conserved.. The lepton charges are omitted in the notations. The $\tau$ leptons are decayed by PYTHIA, and the contribution from the $\tau$ leptonic decay modes leading to two electrons or two muons are included in the signal. Typically, the former final state is obtained with a virtual or real $Z$ boson (Fig. 1 top), and the sensitivity will roughly depend on the mass difference between $H$ and $A$. The latter final state is mediated via a pair of $W$ bosons (Fig. 1 middle) and the sensitivity will instead depend on the difference in mass between $H$ and the charged scalars. The signal is generated fixing the coupling $\lambda_{2}$ to 0.1, $M_{H}$ between 70 and 115 (180) GeV in steps of 5 GeV for $\sqrt{s}$ $240\,(365)\,\text{G\text{eV}}$, and $M_{A}-M_{H}$ between 2 GeV and the kinematic limit for on-shell production in steps of 2 to 5 GeV. Three scenarios are considered for setting $M_{H^{\pm}}$ and $\lambda_{345}$, as detailed in section 2, and 500,000 events are generated per point. The cross section values are shown in Figs. 3 (4) in the simulated $M_{A}-M_{H}$ vs. $M_{H}$ grid plane for $\sqrt{s}=$240 (365) GeV, on the top (left) for the $HH\ell\ell$ final state and on the bottom (right) for the $HH\ell\ell\nu\nu$ final state, for the scenario S1. For the number of events generated, the Monte Carlo (MC) integration error for the cross sections is typically smaller than 0.1%. The minimum transverse momentum of the leptons, $p_{\mathrm{T}}^{\ell}$, is set at 0.5 GeV.
Refer to captionRefer to caption
Figure 3: Cross section from the MG5_aMC@NLO simulation at $\sqrt{s}=$240 GeV, for the points simulated in the $M_{A}-M_{H}$ vs $M_{H}$ plane in scenario S1. All cross sections are in fb, computed for $p_{\mathrm{T}}^{\ell}>0.5$ GeV and with a numerical integration error smaller than 0.1%. Top: $HH\ell\ell$ final state, bottom: $HH\ell\ell\nu\nu$ final state. The dashed line shows the kinematic limit for on-shell production.
@@
Almost no difference is found between S2 and S1, which is expected from the range of $M_{H}$ values of interest here, $M_{H}$$M_{h}$/2/2, leading to the suppression of the BSM decay of $h$ into $HH$. The largest differences are for S3 in the $HH\ell\ell\nu\nu$ final state, for which $M_{H^{\pm}}$ drives the cross section. The cross section ratios between S3 and S2 are shown in Fig. 5 in the $M_{A}-M_{H}$ vs. $M_{H}$ plane for $\sqrt{s}=$240 GeV (left) and $\sqrt{s}=$365 GeV (right) for the $HH\ell\ell\nu\nu$ final state.
@@
Background processes include diboson production ($WW$ and $ZZ$), inclusive $ee$, $\mu\mu$ and $\tau\tau$ production, SM Higgs boson production in association with a $Z$ boson, and $\mathrm{t}\bar{\mathrm{t}}$ production at $\sqrt{s}=$365 GeV. Samples were produced centrally with high statistics, within the so-called Winter2023 campaign [36], as given in Table 2 for $\sqrt{s}=$240 GeV and in Table 3 for $\sqrt{s}=$365 GeV, using PYTHIA v8.306 [37] for the diboson and $\mathrm{t}\bar{\mathrm{t}}$ samples, and WHIZARD v3.0.3 [38] interfaced with PYTHIA v6.4.28 [39] for the other processes. For the inclusive $ee$ production, a lower limit of 30 GeV is set on the invariant mass of the pair of electrons produced. The dielectron channel is hence included in the analysis only in the parameter space where the background is simulated, namely for $M_{A}-M_{H}$>30>30 GeV.
The current target values for the total integrated luminosities of the $\sqrt{s}=$240 and 365 GeV FCC-ee runs are 10.8 and 2.7  ab-1, respectively [40].
Both signal and background samples are processed through DELPHES v3.5.1pre05 [41] with the IDEA [42] detector configuration as per the Winter2023 campaign. Events are produced using the KEY4HEP [43] framework and analysed using the FCCAnalyses package [44].
@@
ZZZZ 56162093 1.359 41
WWWW 373375386 16.4385 23
ZZ(eeee)$h$ 1200000 0.00716 168
ZZ($\mu\mu$)$h$ 1200000 0.00676 178
ZZ($\tau\tau$)$h$ 1200000 0.00675 178
ZZ($\nu\nu$)$h$ 3500000 0.0462 76
ZZ($\mathrm{q}\bar{\mathrm{q}}$)$h$ 6700000 0.136 49
eeee $30<M_{ee}<150$GeV 85400000 8.305 10
@@
Table 2: Background processes generated in the Winter2023 campaign at $\sqrt{s}=$240 GeV. ZZZZ and $WW$ are generated using Pythia8, whilst the others are generated with Whizard+Pythia6. The last column gives the total integrated luminosity simulated.
@@
ZZZZ 11470944 0.6428 18
WWWW 11754213 10.7165 1.1
ZZ(eeee)$h$ 1000000 0.00739 135
ZZ($\mu\mu$)$h$ 1200000 0.004185 287
ZZ($\tau\tau$)$h$ 1100000 0.004172 264
ZZ($\nu\nu$)$h$ 2200000 0.05394 41
ZZ($\mathrm{q}\bar{\mathrm{q}}$)$h$ 2400000 0.032997 73
eeee $30<M_{ee}<150$GeV 3000000 1.527 2.0
@@
$\mathrm{t}\bar{\mathrm{t}}$ 2700000 0.8 3.4
Table 3: Background processes generated in the Winter2023 campaign at $\sqrt{s}=$365 GeV. ZZZZ, WWWW and $t\bar{t}$ are generated using Pythia8, whilst the others are generated with Whizard+Pythia6. The last column gives the total integrated luminosity simulated.
@@ 4 Definition of the objects in Delphes
$\sigma_{E}=\sqrt{E^{2}\times c^{2}+E\times s^{2}+n^{2}}$
@@
Jets are reclustered from Delphes reconstructed particles after removing the selected electron and muon candidates. The Durham algorithm is used, with exclusive clustering of two jets, and an energy-based scheme [45], which leads to good resolution in low-multiplicity final states, and is a standard choice for lepton colliders. The missing energy is taken as given by Delphes, where the DM candidates count as invisible particles.
@@
Tracking $e$, $\mu$, charged hadrons $p_{\text{T}}$ >0.1>0.1 GeV, $|\eta|<2.56$ 1
@@ 5 Background rejection
To remain fully sensitive to all signal points but significantly reduce the background contributions, the following preselection criteria are applied, which are inspired from previous studies performed at generator level with the CLIC setup [12, 13].
Only events with exactly two same-flavour leptons $\ell=e$ or $\mu$ (including decays of $\tau$ leptons to $e$ or $\mu$), $M_{\ell\ell}<120$ GeV, $|p_{z}(\ell\ell)|<70$ GeV are kept. For the $\sqrt{s}=$365 GeV setup, for several benchmark points the signal contribution has $M_{\ell\ell}$ values well above the Z invariant mass. In order to more efficiently reject the backgrounds, the selection is modified to $|p_{z}(\ell\ell)|<140$ GeV, and the selection in $M_{\ell\ell}$ is chosen to be dependent on $p_{z}(\ell\ell)$, with the formula: $M_{\ell\ell}<(-9.0/14.0\times|p_{z}(\ell\ell)|+200)$, with $M_{\ell\ell}$ and $p_{z}(\ell\ell)$ in GeV. As an example, the distribution of $M_{\mu\mu}$ as a function of $p_{z}(\mu\mu)$ is shown for the sum of the major backgrounds and a few signal representative benchmark points from Ref. [12] in Fig. 6, for $\sqrt{s}=$240 (365) GeV on the top (bottom). A good agreement is found with the results obtained using the CLIC setup.
Refer to captionRefer to caption
Figure 6: Dimuon mass $M_{\mu\mu}$ as a function of the dimuon $p_{z}(\mu\mu)$, for the selection of a muon pair with $e^{+}e^{-}$ collisions at $\sqrt{s}=$240 (365) GeV on the top (bottom). The sum of the background processes is shown in red. Several representative signal benchmark points are shown in green (BP1), blue (BP2), yellow (BP7) and pink (BP9). On the bottom, also BP21 is shown in cyan. The specification of the BPs can be found in Ref. [12]. The pink lines highlight the selection requirements.
With a negligible impact on the signal efficiency, the $ee\rightarrow ee,\,\mu\mu,\,\tau\tau$ background is further reduced by 90% if requiring the transverse missing energy $E_{\text{T}}^{miss}$ >5>5 GeV. As no additional object is expected for signal events, a veto is applied on events with any additional electron or muon, any photon with $E>5\,\text{G\text{eV}}$ or any jet. To further reject background, only events with the leading (subleading) lepton $p_{\text{T}}$ <80<80 (60) GeV are kept for $\sqrt{s}=$240 GeV, $p_{\text{T}}$ <140<140 (80) GeV for $\sqrt{s}=$365 GeV. Finally, events with $p_{\ell\ell}/E_{\ell\ell}>0.1$ are selected. All selection criteria are summarised in Table 5.
@@
Preselection $|p_{z}(\ell\ell)|<70$ GeV $|p_{z}(\ell\ell)|<140$ GeV ZZZZ, $ee\rightarrow ee,\,\mu\mu$
$M_{\ell\ell}<120$ GeV $M_{\ell\ell}<(-9.0/14.0\times|p_{z}(\ell\ell)|+200)$ WWWW, $ee\rightarrow ee,\,\mu\mu$
$E_{\text{T}}^{miss}$ >5>5 GeV ZZZZ, $ee\rightarrow ee,\,\mu\mu$
Object veto 3rd lepton $p>5\,\text{G\text{eV}}$     , jet, photon $E>5$ GeV WWWW, ZZZZ, $ee\rightarrow ee,\,\mu\mu,\,\tau\tau$
Leptons $p_{\text{T}}$ $p_{\text{T}}$ <80,60<80,60 GeV $p_{\text{T}}$ <140,80<140,80 GeV WWWW, ZZZZ, $ee\rightarrow\tau\tau$
E/p $p_{\ell\ell}/E_{\ell\ell}>0.1$ $ee\rightarrow ee,\,\mu\mu,\,\tau\tau$
@@
Just from requiring a pair of same-flavour leptons, 16% (4%) of ZZZZ (WWWW) events are selected, further reduced to 5% (2%) after preselection, and 1.5% (1%) after the veto on other objects and final set of selection criteria. For the dilepton production, the preselection achieves 99% (94%) rejection of $ee$ events in the $ee$ channel for $\sqrt{s}=$240 (365) GeV, 99% for the $\mu\mu$ in $\mu\mu$ channels, and 97% for the $\tau\tau$ events in both channels. The object veto and other selection criteria reduce the $ee$, $\mu\mu$ and $\tau\tau$ backgrounds by another factor of about 40%, 70% and 35%, respectively. For the signal, after requiring a pair of same-flavour leptons with p>5p>5GeV  only 2% (0.1%) of events are selected at low mass splitting $M_{A}-M_{H}$, increasing to about 65% (6%) at higher mass splitting, for the production of $HH\ell\ell$ ($HH\ell\ell\nu\nu$) in the scenario S1.
@@
Process eeee $\sqrt{s}=$240 GeV $\mu\mu$ $\sqrt{s}=$240 GeV eeee $\sqrt{s}=$365 GeV $\mu\mu$ $\sqrt{s}=$365 GeV
@@
ZZZZ 9.74e+04 $\pm$ 1.60e+02 1.09e+05 $\pm$ 1.69e+02 9.30e+03 $\pm$ 3.75e+01 1.05e+04 $\pm$ 3.98e+01
WWWW 8.02e+05 $\pm$ 6.18e+02 8.72e+05 $\pm$ 6.44e+02 8.75e+04 $\pm$ 4.64e+02 9.32e+04 $\pm$ 4.79e+02
$\mathrm{t}\bar{\mathrm{t}}$ - - 0 0
eeee 6.16e+05 $\pm$ 8.05e+02 0 1.34e+05 $\pm$ 4.30e+02 0
@@
ZZ(eeee)$h$ 1.11e+02 $\pm$ 2.68e+00 0 1.88e+01 $\pm$ 6.13e-01 9.98e-02 $\pm$ 4.46e-02
ZZ($\mu\mu$)$h$ 6.09e-02 $\pm$ 6.09e-02 1.33e+02 $\pm$ 2.85e+00 9.42e-03 $\pm$ 9.42e-03 1.74e+01 $\pm$ 4.05e-01
ZZ($\nu\nu$)$h$ 2.15e+03 $\pm$ 1.75e+01 2.25e+03 $\pm$ 1.79e+01 6.08e+02 $\pm$ 6.34e+00 6.72e+02 $\pm$ 6.67e+00
ZZ($\tau\tau$)$h$ 1.62e+01 $\pm$ 9.93e-01 1.90e+01 $\pm$ 1.08e+00 2.41e+00 $\pm$ 1.57e-01 2.11e+00 $\pm$ 1.47e-01
ZZ($\mathrm{q}\bar{\mathrm{q}}$)$h$ 0 0 0 0
@@
100-104 $HH\ell\ell$ - - 1.09e+02 $\pm$ 1.59e+00 1.39e+02 $\pm$ 1.80e+00
100-104 $HH\ell\ell\nu\nu$ - - 1.96e+01 $\pm$ 1.00e+00 1.96e+01 $\pm$ 1.00e+00
120-231 $HH\ell\ell$ - - 1.25e+02 $\pm$ 3.47e-01 1.41e+02 $\pm$ 3.68e-01
120-231 $HH\ell\ell\nu\nu$ - - 2.61e-04 $\pm$ 7.59e-06 2.07e-04 $\pm$ 6.77e-06
125-215 $HH\ell\ell$ - - 3.14e+02 $\pm$ 8.65e-01 3.53e+02 $\pm$ 9.18e-01
125-215 $HH\ell\ell\nu\nu$ - - 1.23e-02 $\pm$ 1.14e-04 9.92e-03 $\pm$ 1.03e-04
140-200 $HH\ell\ell$ - - 3.42e+02 $\pm$ 9.29e-01 3.80e+02 $\pm$ 9.79e-01
140-200 $HH\ell\ell\nu\nu$ - - 1.31e-02 $\pm$ 1.21e-04 1.06e-02 $\pm$ 1.09e-04
160-185 $HH\ell\ell$ - - 2.19e+02 $\pm$ 6.29e-01 2.39e+02 $\pm$ 6.56e-01
160-185 $HH\ell\ell\nu\nu$ - - 3.61e-03 $\pm$ 4.81e-05 2.83e-03 $\pm$ 4.26e-05
Table 6: Expected number of events after selection for both channels (eeee and $\mu\mu$) normalised to the total integrated luminosity values of the FCC-ee runs at $\sqrt{s}=$240 and 365 GeV, namely 10.8 and 2.7  ab-1. The signal points are all shown for scenario S1, labelled using the simulated values of $M_{H}$ and $M_{A}$ given in GeV in that order, and separately for the two generated final states, $HH\ell\ell$ and $HH\ell\ell\nu\nu$. Different representative mass points are chosen for the different centre-of-mass scenarios. Only statistical uncertainties are included.